# Sequential Attention-based Sampling for Histopathological Analysis

### v2

@@ 3.3 Targeted State Updater (TSU)
The challenge then is how to effectively integrate this newly acquired high-resolution information into the overall state representation $S_{t}$ without corrupting previously acquired information. One possible approach (e.g., Zhao et al. [36]) involves concatenating $V(a_{t})$ with every patch in $S_{t}$, training an MLP (multi-layer perceptron) with this augmented state as input to produce an updated global state representation $S_{t+1}$ that gradually converges toward the feature representation obtained from all high-resolution patches. We argue that such a global state update is far from ideal. Because only one high resolution patch is observed in each time step, updating instances globally risks corrupting the state, especially when instances unrelated to the sampled patch are concurrently updated. Moreover, information from the sampled patch is likely to be relevant, primarily, for updating instances that share features similar to it [29, 20].
To exploit these correlations, we propose a targeted, similarity-based state update (TSU) model (Fig. 2b). For the selected patch $a_{t}$, we compute the cosine similarity between $S_{t}(a_{t})$ and all other patch representations in $S_{t}$. We then define a set $C$ containing the indices of patches whose cosine similarity with $S_{t}(a_{t})$ exceeds a threshold ($\tau$, a hyperparameter), excluding $a_{t}$ itself. Only the patches in $C$ are updated: $V(a_{t})$ is concatenated with the representations of patches in $C$ and passed through an MLP trained to minimize the MSE loss against the corresponding entries in $V(C)$. Additionally, $S_{t}(a_{t})$ is directly replaced with $V(a_{t})$ because its exact high-resolution representation is available. In the next time step, the patch $a_{t}$ is masked out, and future actions are sampled from the remaining unmasked patches; this strategy is applied during both the training and evaluation stages. With ablation studies (Table 2), we show that this selective update strategy outperforms naive approaches.
@@
  Input: Current state $S_{t}\in\mathbb{R}^{N\times d}$; Set $I$ of sampled patch indices, upto time $t$
@@
1:  $a_{t}\sim\pi_{\theta_{t}}(.\mid S_{t})$Sample patch index $a_{t}$ from the policy distribution}
2:  $I\leftarrow I\cup\{a_{t}\}$Store selected patch index}
@@
5:     Update $\pi_{\theta}$ with a PPO algorithm; intermediate reward $r_{t}=-BCE(y,\hat{y_{t}})$
@@
7:  $V(a_{t})\leftarrow f_{H}(U_{a_{t}};\theta_{H})$extract features $V(a_{t})\in\mathbb{R}^{d}$ from high-resolution zoomed in features, $U_{a_{t}}\mathbb{R}^{k\times d}$ of the sampled patch, with the HAFED model}
8:  $S_{t+1}(a_{t})\leftarrow V(a_{t})$update the sampled patch with its high-resolution feature representation}
9:  $C\leftarrow\{i:\cos\angle(S_{t}(i),S_{t}(a_{t}))\geq\tau\}$identify set of patch indices whose cosine similarity with $S_{t}(a_{t})$ exceeds a threshold}
@@ 3.4 Sequential Attention-based Sampling for Histopathological Analysis
In this section, we describe how we train a reinforcement learning (RL) agent to select which patches to zoom into, leveraging feature distillation with the HAFED model and state update with the TSU model. We formulate the patch selection process as a sequential decision-making problem, where an agent interacts with a slide-specific environment over multiple steps (Algorithm 1).
During each training episode, at each timestep $t$, the agent observes the WSI state $S_{t}$, which is initialized with the low-resolution feature $S_{0}=Z$. Based on $S_{t}$, the policy network $\pi_{\theta}$ produces a probability distribution over a set of actions from a discrete action space $\{1,2,\ldots,N\}$. The agent samples an action $a_{t}$, corresponding to the next low-resolution patch to zoom in, according to the current policy $\pi_{\theta_{t}}(\cdot\mid S_{t})$ with a stochastic sampling rule. The zoomed-in patch is passed through HAFED to obtain its high-resolution representation $V(a_{t})$. This representation is then used by the TSU to selectively refine the current state $S_{t}$, by modifying only those patches that are similar to $a_{t}$ in the feature space. This produces an updated state $S_{t+1}$. The classifier trained together with HAFED then uses the updated state $S_{t+1}$ to produce a slide-level prediction $\hat{y}_{t}$. This prediction is compared against the ground truth label $y$ to compute a time-step wise reward, based on a negative binary cross-entropy (BCE) loss. This feedback is used to update the policy network using Proximal Policy Optimization (PPO) [28] algorithm. We employed Generalized Advantage Estimation (GAE) in computing advantage function [27], whose output was trained to predict the reward until the end of the episode. Both the policy and value network were modeled on standard architectures [14] details are in Appendix A.2.3. The same sequence of steps was followed during evaluation, except that the action with the highest probability based on the policy network’s output was chosen deterministically.
@@ 4.1 Dataset and implementation details
We evaluate the effectiveness of SASHA with two publicly available benchmark datasets CAMELYON16 breast cancer [3] and TCGA-NSCLC lung cancer dataset [23] dataset. While the CAMELYON16 labels correspond to normal versus pathological WSIs, TCGA-NSCLC labels correspond to distinguishing two different types of cancers, adenocarcinoma versus squamous cell carcinoma. We followed standardized train-test splits for both datasets. Further details on the datasets, hyperparameter selection, implementation, and training are provided in Appendix A.1 and A.2.
@@
Sample 100% Max Pooling 0.933 $\pm$0.007 0.971$\pm$0.001 0.911$\pm$0.008 0.881$\pm$0.010 0.952$\pm$0.008 0.882$\pm$0.010
ABMIL 0.947$\pm$0.008 0.958$\pm$0.016 0.928$\pm$0.011 0.883$\pm$0.045 0.947$\pm$0.021 0.884$\pm$0.045
CLAM_SB 0.949$\pm$0.012 0.976$\pm$0.008 0.931$\pm$0.015 0.905$\pm$0.019 0.961$\pm$0.009 0.907$\pm$0.018
TransMIL 0.936$\pm$0.006 0.953$\pm$0.008 0.912$\pm$0.008 0.898$\pm$0.015 0.964$\pm$0.004 0.899$\pm$0.016
DTFD 0.944$\pm$0.014 0.977$\pm$0.008 0.944$\pm$0.014 0.911$\pm$0.012 0.962$\pm$0.010 0.911$\pm$0.022
ACMIL 0.941$\pm$0.015 0.970$\pm$0.011 0.924$\pm$0.017 0.906$\pm$0.025 0.959$\pm$0.006 0.907$\pm$0.026
HAFED (ours) 0.963$\pm$0.008 0.980$\pm$0.003 0.951$\pm$0.011 0.923$\pm$0.011 0.966$\pm$0.015 0.925$\pm$0.010
10% RLogist-0.1 0.824 0.829 0.828 0.892
SASHA-0.1 (ours) 0.901$\pm$0.021 0.918$\pm$0.014 0.856$\pm$0.031 0.897$\pm$0.023 0.956$\pm$0.023 0.898$\pm$0.024
20% RLogist-0.2 0.862 0.879 0.839 0.903
SASHA-0.2 (ours) 0.953$\pm$0.017 0.979$\pm$0.008 0.937$\pm$0.024 0.912$\pm$0.010 0.963$\pm$0.014 0.914$\pm$0.011
Table 1: Performance comparison on the CAMELYON16 and TCGA-NSCLC datasets. The top 7 rows reflect models that process all patches at high resolution (“Sample 100%”). Underlined values indicate the best performance among these methods. Rows 8-9 and 10-11 reflect models that selectively sample either “10%” or “20%” of the patches, respectively. Bold values indicate the best performance among these methods at the respective sampling fraction.
@@ 4.2 Baselines and Metrics
We evaluate and compare the performance of our model against two types of baselines. First, we evaluate a range of different sota methods that analyze all patches in the WSI at high resolution, including MaxPooling, ABMIL [14], CLAM [20], TransMIL [29], DTFD [34], and ACMIL [35]. We include also comparisons with our hierarchical attention-based (HAFED) model (Section 3.2) with the full resolution WSI. Yet, these methods serves only as an upper bound on achievable performance.
For a fairer comparison, we evaluate the performance of models that sample only a fraction of the WSI. Specifically, we evaluate RLogist trained to sample either exactly $10\%$ or exactly $20\%$ of patches at high resolution and compare it with our model trained under identical settings; this analysis also enables analyzing the impact of observation budgets on model performance. In addition, we also compare two other sota attention-based models ACMIL and DTFD, but evaluate their performance under same set of observations budgets, with random sampling of high-resolution patches (Table 2); details in Appendix A.3. In all cases, models were trained with the official code provided by the respective studies, except for RLogist. The official RLogist codebase lacks major elements, including state update models, possibly due to proprietarity; in this case, we report the performance metrics as reported in the original study on the same benchmarks [36].
Performance metrics are reported in terms of Accuracy, AUC, and F1 scores for normal versus pathological tissue classification (CAMELYON16) or for classifying between the different tumor types (TCGA-NSCLC). In addition, we expect that selective sampling models, such as ours, would exhibit significantly faster inference times compared to models that examine all patches at high resolution; therefore, we also evaluate and compare inference times (Fig. 3c) . Finally, we also compute a metric of “WSI compressibility”. We define this as a ratio of the original WSI image size, at full resolution, to the size the of WSI representation used by each model’s classifier to make a final decision. This metric has important real-world implications in terms of efficient archival and storage of WSI images. We compare models of comparable complexity and performance (CLAM, TransMIL, ACMIL, DTFD, SASHA) in terms both WSI compressibility and overall model size.
@@ 4.3 Classification Performance, Inference Time and Compressibility
Classification Performance. SASHA consistently outperforms the sota competing deep RL model – RLogist – by a large margin, for both the 10% and 20% observation budgets (Table 1, SASHA-0.1 and SASHA-0.2, vs RLogist-0.1 and RLogist-0.2, respectively). Improved classification performance was reflected in all 3 metrics quantified: Accuracy, AUC and F1-score. Moreover, our HAFED model performed on par with the best sota models (e.g. ACMIL, DTFD) that analyze the full, high-resolution WSI (Table 1, underlined). Interestingly, even with sampling 5x fewer patches, SASHA-0.2 nears ceiling performance achievable with HAFED, when all 100% of patches are sampled at high resolution (Table 1, compare SASHA-0.2 with HAFED).
Inference time. The average inference time per WSI for the CAMELYON16 dataset, including both pre-processing and feature extraction time, was $\sim$117 seconds, using HAFED (Fig. 3c). Other sota attention-based methods like ACMIL and DTFD also exhibited similar average inference times per WSI. The most expensive operation was feature extraction for each of the high resolution patches. By contrast, our model samples only a few patches at high-resolution, thereby leading to a substantial decrease in overall inference time, $\sim$14 seconds and $\sim$26 seconds per WSI for SASHA-0.1 and SASHA-0.2 – an 8$\times$ and 4$\times$ speedup, respectively – compared to the second fastest model (Fig. 3c).
WSI compressibility. Similarly, a comparison of the WSI compressibility revealed a significant (>16×>16$\times$> 16 ×) improvement, compared to models that examine the entire slide at high resolution (Fig. 3d). The main reason for this saving is the fact that our HAFED model encodes features for both low- and high-resolution patches in a uniform $N\times d$ dimensional space; by contrast, other models encode high-resolution features in a $N\times k\times d$ dimensional space, where $k$ reflects the ratio of the zoomed-in to the scanning resolution size. The full set of results, including a comparison of model sizes, is shown in Appendix A.4. These results highlight the efficiency of our method in balancing performance with computational and representational costs.
@@ 4.4 Ablation Study and Limited Sampling
We performed a systematic ablation study removing each component of the model in turn – or replacing it with a naive variant – and evaluating it with the CAMELYON16 dataset. These included: a) replacing the WSI-pretrained ViT with a ResNet50, used commonly for feature extraction [37, 36], b) replacing multi-branch attention with single branch in the Classifer [14], c) replacing the targeted state update policy with a global state update [36], d) removing the TSU and updating the state only of the sampled (local, high-resolution) patch, and e) and selecting an action based on a random policy (see “Variant”s in Table 2). In every case, we observed sharp drops in accuracy, ranging from 7-45%, relative to the baseline SASHA model; similar drops occurred in AUROC and F1 scores Table 2. The strongest impact occurred upon changing the RL policy (Table 2, e), with the next most impactful effects occurring upon changing the feature extractor (a) or the state update method (c-d). The least impact occurred upon changing the classifier to single branch attention (b), but even here, the F1 score dropped by 14% points. Additional ablation experiments, and including exploring the effect of terminal versus intermediate reward during RL agent training, adopting a stochastic action selection policy are described in Appendix A.3. These results indicate that every component of our model was critical for achieving high performance, comparable to sota models.
Finally, we examined the effect of imposing a limited observation budget on other sota models, and compared the performance of SASHA under the same budget. Specifically, ACMIL and DTFD were trained and tested with a random selection of high-resolution patches under a 20% observation budget for each WSI. In this case also we observed substantial drops in performance relative to the default SASHA model (Table 2, f-g). In other words, randomly sampling a limited fraction of patches at high resolution, even with the best sota models does not compare with SASHA’s performance with sampling the same, respective fraction.
@@
(b) 1o Att. Branch * 0.899 0.964 0.851
(c) Global Update * 0.837 0.824 0.779
(d) Local Update * 0.860 0.901 0.800
(e) Random Policy * 0.516 0.550 0.553
(f) ACMIL-0.2 (Rnd.) * * 0.838 0.857 0.773
(g) DTFD-0.2 (Rnd.) * * 0.862 0.878 0.862
@@ 4.5 Patch selection strategy: Explainability
When deploying deep RL models in medical settings it is essential that they learn not only an effective, but also an explainable action policy. For this, we explore whether SASHA’s RL agent exhibits a meaningful patch selection strategy prior to classifying WSI images (Fig. 3a and Fig. 3b). We propose that such a strategy must satisfy at least two important requirements: i) In WSIs containing tumor tissue, the RL agent must preferentially sample patches with a large fraction of tumor tissue for high-resolution scans, and ii) an effective RL policy should preferentially sample patches afforded a high attention score by sota models that examine the entire WSI at high resolution. We evaluated each of these criteria in the data. Indeed, the tumor fraction in patches sampled by the RL agent at high resolution (10% with SASHA-0.1 or 20% with SASHA-0.2) significantly exceeds that in a corresponding number of randomly chosen patches, among those not sampled by the agent (Fig. 3e, p<0.001). Moreover, the RL policy guides the agent to preferentially samples patches afforded a high attention score by sota models: both the average attention score (Fig. 3g, p<0.001) and the fraction of top-k patches sampled under the RL policy were statistically significantly greater than the fraction of patches sampled by a random policy (Fig. 3f, p<0.001). In other words, our RL agent’s sampling actions were guided by an efficient, intuitive and explainable policy.

### v3

@@ 3.3 Targeted State Updater (TSU)
The challenge then is how to effectively integrate this newly acquired high-resolution information into the overall state representation $S_{t}$ without corrupting previously acquired information. One possible approach (e.g., zhao2022rlogist) involves concatenating $V(a_{t})$ with every patch in $S_{t}$, training an MLP (multi-layer perceptron) with this augmented state as input to produce an updated global state representation $S_{t+1}$ that gradually converges toward the feature representation obtained from all high-resolution patches. We argue that such a global state update is far from ideal. Because only one high resolution patch is observed in each time step, updating instances globally risks corrupting the state, especially when instances unrelated to the sampled patch are concurrently updated. Moreover, information from the sampled patch is likely to be relevant, primarily, for updating instances that share features similar to it [shao2021transmil, lu2021clam].
To exploit these correlations, we propose a targeted, similarity-based state update (TSU) model (Fig.˜2b). For the selected patch $a_{t}$, we compute the cosine similarity between $S_{t}(a_{t})$ and all other patch representations in $S_{t}$. We then define a set $C$ containing the indices of patches whose cosine similarity with $S_{t}(a_{t})$ exceeds a threshold ($\tau$, a hyperparameter), excluding $a_{t}$ itself. Only the patches in $C$ are updated: $V(a_{t})$ is concatenated with the representations of patches in $C$ and passed through an MLP trained to minimize the MSE loss against the corresponding entries in $V(C)$. Additionally, $S_{t}(a_{t})$ is directly replaced with $V(a_{t})$ because its exact high-resolution representation is available. In the next time step, the patch $a_{t}$ is masked out, and future actions are sampled from the remaining unmasked patches; this strategy is applied during both the training and evaluation stages. With ablation studies (Table˜2), we show that this selective update strategy outperforms naive approaches.
@@
Input: Current state $S_{t}\in\mathbb{R}^{N\times d}$; Set II of sampled patch indices, upto time $t$
@@
[…]
@@
[…]
@@
[…]
@@ 3.4 Sequential Attention-based Sampling for Histopathological Analysis
We train a reinforcement learning (RL) agent to select which patches to zoom into, leveraging feature distillation with the HAFED model and state update with the TSU model. We formulate the patch selection process as a sequential decision-making problem, where an agent interacts with a slide-specific environment over multiple steps (Algorithm˜1). During each training episode, at each timestep $t$, the agent observes the WSI state $S_{t}$, which is initialized with the low-resolution feature $S_{0}=Z$. Based on $S_{t}$, the policy network $\pi_{\theta}$ produces a probability distribution over a set of actions from a discrete action space $\{1,2,\ldots,N\}$. The agent samples an action $a_{t}$, corresponding to the next low-resolution patch to zoom in, according to the current policy $\pi_{\theta_{t}}(\cdot\mid S_{t})$ with a stochastic sampling rule. The zoomed-in patch is passed through HAFED to obtain its high-resolution representation $V(a_{t})$. This representation is then used by the TSU to selectively refine the current state $S_{t}$, by modifying only those patches that are similar to $a_{t}$ in the feature space. This produces an updated state $S_{t+1}$. The classifier trained together with HAFED then uses the updated state $S_{t+1}$ to produce a slide-level prediction $\hat{y}_{t}$. This prediction is compared against the ground truth label $y$ to compute a time-step-wise reward, based on a negative cross-entropy (CE) loss. This feedback is used to update the policy network using Proximal Policy Optimization (PPO) [schulman2017ppo] algorithm. We employed Generalized Advantage Estimation (GAE) in computing the advantage function [gae], whose output was trained to predict the reward until the end of the episode. Both the policy and value network were modeled on standard architectures [ilse2018abmil] (details in Appendix A.2.3). The same steps were followed during evaluation, except that the action with the highest probability based on the policy network’s output was chosen deterministically.
@@ 4.1 Dataset and implementation details
We evaluate the effectiveness of SASHA with two publicly available benchmark datasets CAMELYON16 breast cancer [bejnordi2017diagnostic] and TCGA-NSCLC lung cancer dataset [tcga2014cancer]. While the CAMELYON16 labels correspond to normal versus pathological WSIs, TCGA-NSCLC labels correspond to distinguishing two different types of cancers, adenocarcinoma versus squamous cell carcinoma. We followed standardized train-test splits for both datasets. In addition, we also evaluate SASHA on multi-class classification tasks using the BRACS and Camelyon+ datasets. Further details on the datasets, hyperparameter selection, implementation, and training are in Appendices A.1 and A.2.
@@
Sample 100%
[…]
k
Zoom-MIL (k=300) 0.780 $\pm$0.017 0.700 $\pm$0.029 0.738 $\pm$0.015 0.908 $\pm$0.004 0.967 $\pm$0.004 0.908 $\pm$0.004
10%
RLogist-0.1 0.824 0.829 0.828 0.892
[…]
20%
RLogist-0.2 0.862 0.879 0.839 0.903
[…]
Table 1: Performance comparison on the CAMELYON16 and TCGA-NSCLC datasets. The top 7 rows reflect models that process all patches at high resolution (“Sample 100%”). Underlined values indicate the best performance among these methods. Rows 9-11 and 12-14 reflect models that selectively sample either “10%” or “20%” of the patches, respectively. Bold values indicate the best performance among these methods at the respective sampling fraction.
@@ 4.2 Baselines and Metrics
We evaluate and compare the performance of our model against two types of baselines. First, we evaluate a range of different sota methods that analyze all patches in the WSI at high resolution, including MaxPooling, ABMIL [ilse2018abmil], CLAM [lu2021clam], TransMIL [shao2021transmil], DTFD [zhang2022dtfd], and ACMIL [zhang2023acmil]. We also include comparisons with our hierarchical attention-based (HAFED) model (Section˜3.2) with the full-resolution WSI. Yet, these methods serve only as an upper bound on achievable performance.
For a fair comparison, we also evaluate models that sample only a fraction of the WSI. Specifically, we compare our approach with RLogist [zhao2022rlogist] when trained to sample exactly $10\%$ and $20\%$ of patches at high resolution, under identical experimental settings. We also compare with a non-RL-based Zoom-MIL [zoom_mil] model that samples exactly k = 80, 160, or 300 patches at low resolution (5x $\rightarrow$ 20x). The choice of k values aligns with the patch statistics of the dataset: WSIs contain $\sim$800 low-resolution patches on average; thus, $k=80$ and $k=160$ correspond to $10\%$ and $20\%$ of the available patches, on average, providing a fair comparison with the RLogist-0.1/SASHA-0.1 and RLogist-0.2/SASHA-0.2 settings, respectively. $k=300$ follows the default configuration used in the original Zoom-MIL paper [zoom_mil]. This setup allows us to systematically analyze the effect of varying observation budgets on model performance. In addition, we also compare against ACMIL and DTFD under the same observation budgets, with random sampling of high-resolution patches (Table˜2); details in Appendix A.3. Models were trained with the official codebase provided by the respective studies, except for RLogist. The RLogist codebase is insufficient for replication; therefore, we report the performance metrics as is from the original study [zhao2022rlogist].
Our performance metrics include Accuracy, AUC, and F1 scores for normal versus pathological tissue classification (CAMELYON16) or for classifying between the different tumor types (TCGA-NSCLC). In addition, we expect that selective sampling models, such as ours, would exhibit significantly faster inference times compared to models that examine all patches at high resolution; therefore, we also evaluate and compare inference times (Fig.˜3c) . Finally, we also compute a metric of “WSI compressibility”. We define this as a ratio of the original WSI image size, at full resolution, to the size the of WSI representation used by each model’s classifier to make a final decision. This metric has important real-world implications in terms of efficient archival and storage of WSI images. We compare models of comparable complexity and performance (CLAM, TransMIL, ACMIL, DTFD, SASHA) in terms both WSI compressibility and overall model size.
@@ 4.3 Classification Performance, Inference Time and Compressibility
Classification Performance. SASHA consistently outperforms the sota competing deep RL model – RLogist – by a large margin, for both the 10% and 20% observation budgets (Table˜1, SASHA-0.1 and SASHA-0.2, vs RLogist-0.1 and RLogist-0.2, respectively). Improved classification performance was reflected in all 3 metrics quantified: Accuracy, AUC, and F1-score. Zoom-MIL performed comparably with SASHA-0.2 for the TCGA benchmark, but underperformed heavily for CAMELYON-16. Moreover, our HAFED model performed on par with the best sota models (e.g. ACMIL, DTFD) that analyze the full, high-resolution WSI (Table˜1, underlined). Interestingly, even with sampling 5x fewer patches, SASHA-0.2 nears ceiling performance achievable with HAFED, when all 100% of patches are sampled at high resolution (Table˜1, compare SASHA-0.2 with HAFED).
[…]
WSI compressibility. Similarly, a comparison of the WSI compressibility revealed a significant ($16\times$) improvement, compared to models that examine the entire slide at high resolution (Fig.˜3d). The main reason for this saving is the fact that our HAFED model encodes features for both low- and high-resolution patches in a uniform $N\times d$ dimensional space; by contrast, other models encode high-resolution features in a $N\times k\times d$ dimensional space, where $k$ reflects the ratio of the zoomed-in to the scanning resolution size. The full set of results is shown in Appendix A.4. These results highlight the efficiency of our method in balancing performance with computational and representational costs.
@@ 4.4 Ablation Study and Limited Sampling
We performed a systematic ablation study removing each component of the model in turn – or replacing it with a naive variant – and evaluating it with the CAMELYON16 dataset. These included: a) replacing the WSI-pretrained ViT with a ResNet50, used commonly for feature extraction [zheng2024cvpr, zhao2022rlogist], b) replacing ViT with the CONCH medical image-encoder[lu2023conch], pretrained on 1.17M image caption pairs, c) replacing multi-branch attention with single branch in the Classifer [ilse2018abmil], d) replacing the targeted state update policy with a global state update [zhao2022rlogist], e) removing the TSU and updating the state only of the sampled (local, high-resolution) patch, and f) and selecting an action based on a random policy (see “Variant”s in Table˜2). In every case, we observed sharp drops in accuracy, ranging from 7-45%, relative to the baseline SASHA model; similar drops occurred in AUROC and F1 scores Table˜2. The strongest impact occurred when changing the RL policy (Table˜2, f). The next strongest impact occurred when changing the feature extractor to ResNet-50 (a), but this impact was mitigated when using the CONCH encoder (b). Intermediate levels of impact occurred when changing the classifier to single-branch attention (c) or altering the state update method (d-e). Additional ablation experiments are described in Appendix A.3. These results indicate that every component of our model was critical for achieving high performance, comparable to sota models.
Finally, we examined the effect of imposing a limited observation budget on other sota models, and compared the performance of SASHA under the same budget. Specifically, ACMIL and DTFD were trained and tested with a random selection of high-resolution patches under a 20% observation budget for each WSI. In this case also we observed substantial drops in performance relative to the default SASHA model (Table˜2, g-h). In other words, randomly sampling a limited fraction of high-resolution patches, even with the best models, failed to outperform SASHA.
@@
(b) CONCH * 0.930 0.950 0.905
(c) 1o Att. Branch * 0.899 0.964 0.851
(d) Global Update * 0.837 0.824 0.779
(e) Local Update * 0.860 0.901 0.800
(f) Random Policy * 0.516 0.550 0.553
(g) ACMIL-0.2 (Rnd.) * * 0.838 0.857 0.773
(h) DTFD-0.2 (Rnd.) * * 0.862 0.878 0.862
@@ 4.5 Patch selection strategy: Explainability
When deploying deep RL models in medical settings, it is essential that they learn not only an effective, but also an explainable action policy. For this, we explore whether SASHA’s RL agent exhibits a meaningful patch selection strategy prior to classifying WSI images (Fig.˜3a and Fig.˜3b). We propose that such a strategy must satisfy at least two important requirements: i) In WSIs containing tumor tissue, the RL agent must preferentially sample patches with a large fraction of tumor tissue for high-resolution scans, and ii) an effective RL policy should preferentially sample patches afforded a high attention score by sota models that examine the entire WSI at high resolution. We evaluated each of these criteria in the data. Indeed, the tumor fraction in patches sampled by the RL agent at high resolution (10% with SASHA-0.1 or 20% with SASHA-0.2) significantly exceeds that in a corresponding number of randomly chosen patches, among those not sampled by the agent (Fig.˜3e, p<0.001). Moreover, the RL policy guides the agent to preferentially samples patches afforded a high attention score by sota models: both the average attention score (Fig.˜3g, p<0.001) and the fraction of top-k patches sampled under the RL policy were statistically significantly greater than the fraction of patches sampled by a random policy (Fig.˜3f, p<0.001). In other words, our RL agent’s sampling actions were guided by an efficient, intuitive and explainable policy.